OpenStudy (anonymous):

Write the equation of the line that passes through (2, 3) and is parallel to the line 12x – 5y = 2.

OpenStudy (nirmalnema):

parallel lines have same slope. its slope is -a/b =12/5 use one point form.. y-3=12/5 (x-2) 5y-15=12x-24 12-5y-9=0 is the required eq

OpenStudy (nirmalnema):

@Bayleefishee you got it